# Sensitivity of CTAO to axion-like particles from blazars: a machine learning approach

### v1

@@ Abstract
Blazars are a class of active galactic nuclei, supermassive black holes located at the centres of distant galaxies characterised by strong emission across the entire electromagnetic spectrum, from radio waves to gamma rays. Their relativistic jets, closely aligned to the line of sight from Earth, are a rich and complex environment, characterised by the presence of strong magnetic fields over parsec-scale lengths. Owing to their cosmological distance from Earth, these sources serve as ideal targets to probe non-standard gamma-ray propagation. In particular, axion-like particles (ALPs) could be detected through their coupling to photons, which enables ALP-photon conversions in external magnetic fields, leading to distinct signatures in the blazars’ gamma-ray spectra. In this work, we explore a novel approach to constrain the ALP parameter space using gamma-ray observations, based on the use of machine-learning classifiers. We apply this technique to simulated observations of two bright blazars – Mrk 501 and PKS 2155-304 – with the Cherenkov Telescope Array Observatory (CTAO), a next-generation gamma-ray facility well suited to probe such features, thanks to its improved energy resolution and point-source sensitivity with respect to present ground-based gamma-ray telescopes. The obtained $2\sigma$ exclusion regions on the ALP parameter space are consistent with those found by applying a standard likelihood-ratio test, and suggest that the CTAO sensitivity to ALPs could be extended beyond existing constraints over a wide mass range.
@@ 1 Introduction
Axion-like particles (ALPs) are a kind of hypothetical, pseudo-scalar elementary particles whose existence is predicted by several extensions of the Standard Model. Their name is derived from the original axion model, which was formulated by Peccei and Quinn [pecceiquinn1977], Weinberg [weinberg1978], and Wilczek [wilczek1978] as a solution of the strong CP problem in the context of quantum chromodynamics (QCD). While the existence of this original axion was soon ruled out [donnelly1978, zehnder1981], several “invisible axion” models were proposed, such as KSVZ [kim1979, shifman1980] and DFSZ [dine1981, zhitnitsky1980], which could elude experimental constraints while still solving the strong CP problem. In addition to these, it became clear that axion-like particles could be a more general prediction of beyond-the-Standard-Model theories such as Grand Unification Theories and string theories. These ALPs are not necessarily related to the original Peccei-Quinn model, but, like the QCD axion, they often arise as pseudo-Nambu-Goldstone bosons of spontaneously broken symmetry groups. Axions and ALPs are currently of great interest in cosmology and astrophysics, as they are among the best-motivated candidates as constituents of dark matter. Reviews of models and experimental searches for both QCD axions and ALPs can be found e.g. in refs. [jaeckel2010, ringwald2012, irastorza2018, diluzio2020].
@@
The most important feature for the experimental detection of ALPs is their effective two-vertex coupling to photons, whose strength is controlled by the coupling constant $g_{a\gamma}$. Such a coupling is the basis for the most common approach to ALP searches, which exploits the expected ALP-photon mixing in the presence of external magnetic fields [sikivie1983, raffelt1988]. This effect is employed, for example, in beam regeneration (or “light shining through walls”, LSW) experiments, such as ALPS-II [bahre2013], where a laser beam is shot against an opaque screen; in the presence of a strong enough magnetic field, some photons in the beam should be able to convert into ALPs, pass through the screen and oscillate back to photons on the other side. Other experiments for the direct detection of ALPs include helioscopes, like CAST [zioutas1999] and the foreseen IAXO [armengaud2019], and resonant cavity experiments such as ADMX [du2018] and MADMAX [brun2019].
@@
As a general property of QCD axions, the axion-photon coupling $g_{a\gamma}$ is proportional to the axion mass $m_{a}$, as both quantities are inversely proportional to the Peccei-Quinn energy scale $f_{a}$ (e.g. [irastorza2018]). On the other hand, in the more general case of ALPs there is typically no fixed relationship between $m_{a}$ and $g_{a\gamma}$, so that they are independent coordinates spanning a 2-dimensional parameter space.
@@
Astrophysical gamma-ray sources characterised by one or more magnetic field environments along their line of sight to Earth offer a natural framework to probe ALP-related effects, which has been extensively discussed in literature [hooper2007, mirizzi2007, mirizzi2008, deangelis2007, deangelis2008, deangelis2011, horns2012, batkovic2021, galanti2022]. Gamma-ray polarization effects could in principle observed, as ALPs only couple to the photon state parallel to the external transverse magnetic field; however, such effects are not detectable with present-day instruments [bassan2010, galanti2024]. On the other hand, two main measurable spectral features may be investigated, whose extent generally depends on the field strength and topology, ALP parameters, source distance, and energy range of interest. The first is the presence of energy-dependent wiggles superimposed on the spectral shape, due to the peculiar shape of the photon survival probability $P_{\gamma\gamma}$ as a function of energy. The second is a spectral hardening at very high energy (VHE), due to gamma-ray photons converting into ALPs in the magnetic field of the source and back into photons in the Milky Way. This mechanism would allow the ALP-photon beam to travel cosmological distances, eluding the absorption that is normally experienced by VHE gamma rays due to the pair production (Breit-Wheeler scattering [breitwheeler1934]) on the soft photons making up the extragalactic background light (EBL) [dominguez2011, franceschini2008, finke2010, gilmore2012, saldanalopez2021, deangelis2013]. The resulting effect, analogous to a “cosmic LSW experiment”, is expected to be more significant in the spectral tails of distant sources, for which the EBL absorption in the absence of ALPs would be strongest.
@@
A typical target of these studies is represented by blazars. These sources are a particular class of active galactic nuclei (AGN), supermassive black holes characterised by accretion disks and strong relativistic jets at a small angle ($\lesssim 20\degree$) with the line of sight from Earth [padovani2017]. Their broadband spectral energy distribution (SED) is characterised by a low-energy component due to the synchrotron radiation emitted by accelerated electrons in the jet, peaked between infrared and X-ray frequencies, and higher-energy emission in the gamma rays that may be produced either by leptonic or hadronic mechanisms [bottcher2013, cerruti2020]. Leptonic mechanisms typically assume the gamma-ray emission to be originated by the same electron population that produces the synchrotron radiation via inverse Compton scattering, either on the synchrotron radiation field itself (synchrotron self-Compton, SSC) or on an external low-energy photon field (such as thermal photons from the broad line region, accretion disk or dusty torus, or the cosmic microwave background). In hadronic models, high-energy photons are emitted by a population of relativistic protons and other nuclei in the jet, typically via synchrotron radiation or photo-meson production, with the associated production of neutrinos. Based on their optical properties, blazars can be classified between BL Lacertae (BL Lac) objects and flat spectrum radio quasars (FSRQs). In addition, they can be distinguished between high-, intermediate- and low-synchrotron peaked (HSP, ISP and LSP) depending on their peak synchrotron frequency. Many blazars display significant spectral variability over timescales ranging from years down to minutes [bottcher2019], with high-flux flaring states representing important opportunities of detection or even discovery of new sources. Despite being a very rare class of objects, almost all extragalactic sources detected in gamma-rays are associated with blazars.
@@
The relativistic jets of blazars typically feature magnetic fields of order $\mathcal{O}(\rm mG)$ to $\mathcal{O}(\rm G)$ over $\mathcal{O}(\rm pc)$ scales. These objects may also be located inside galactic clusters with complex magnetic field structures of $\mathcal{O}(\rm\upmu G)$ intensity over $\mathcal{O}(\rm Mpc)$ scales; a notable example of this environment is represented by the Perseus galaxy cluster, which hosts the Seyfert galaxy NGC 1275. Additionally, being extragalactic sources, blazars are located at a cosmological distance from Earth (represented by the redshift parameter $z$), which makes EBL absorption effects important. All these characteristics may allow for hints of ALP-photon conversions in the gamma-ray spectra of blazars. Several constraints on the ALP parameter space have been obtained by analysing gamma-ray data from different blazars [abramowski2013, jacobsen2023, davies2023] and from the Perseus galaxy cluster [ajello2016, abe2024].
@@
In this context, the next-generation Cherenkov Telescope Array Observatory (CTAO [^1] https://www.ctao.org/, formerly known as CTA) will be an ideal instrument to probe this kind of phenomena. The observatory will be made up of three different types of imaging atmospheric Cherenkov telescopes (IACTs), distributed in arrays over two sites in the Northern and Southern hemispheres [acharya2013]. The Northern Array will be located at Observatorio del Roque de los Muchachos in La Palma, Canary Islands, and will be made up of 4 Large-Sized Telescopes (LSTs) and 9 Medium-Sized Telescopes (MSTs); the Southern Array, in Paranal, Chile, will feature 14 MSTs and 37 Small-Sized Telescopes (SSTs). These different configurations were designed to obtain the best sensitivity to sources visible from the two hemispheres, with the Southern array optimised to detect the highest energy photons (up to several hundreds of TeV) from the Galactic centre. In addition to this baseline “Alpha” configuration, an upgrade is foreseen by the Italian CTA+ Program [^2] https://pnrr.inaf.it/progetto-ctaplus/ to add 2 LSTs [antonelli2023] and 5 more SSTs to the Southern site. An alternative configuration has also been proposed, using telescopes of the Schwarzschild-Couder type [adams2021] to complement the standard MSTs (which feature a Davies-Cotton optical design). CTAO will obtain a significant improvement in all areas of performance with respect to present-days IACT systems: in particular, its greatly enhanced energy resolution and point-source sensitivity will make it an ideal instrument to study the spectra of distant sources in search of potential non-standard spectral features [ctascience2018]. The CTAO sensitivity over the ALP parameter space has been estimated by simulating observations of various blazars [meyer2014a, meyer2014b] and of the Perseus galaxy cluster [abdalla2021].
@@
Almost all of the previously cited works rely on the definition of some kind of test statistic (TS) based on the logarithm of the likelihood ratio (equivalently, the difference in log-likelihoods) to test the ALP hypothesis on real or simulated data [jacobsen2023, davies2023, ajello2016, abe2024, meyer2014a, meyer2014b, abdalla2021]. In mathematical terms, the TS is given by
${\rm TS}=-2\ln\frac{\mathcal{L}(H_{1})}{\mathcal{L}(H_{0})}=-2(\ln\mathcal{L}(H_{1})-\ln\mathcal{L}(H_{0}))\,$, (1.1)
@@ 2.1 Source selection
The observation of AGNs is one of the Key Science Projects (KSPs) proposed for CTAO, to be carried out within the first 10 years of operation of the observatory [ctascience2018]. The programme of this KSP will be organized in three parts: long-term monitoring of known AGNs, search and follow-up of flares, and measurement of high-quality spectra for different AGN classes and redshifts. One of the declared goals of the AGN KSP is to study possible imprints of new fundamental physics on the cosmological propagation of gamma-rays, including ALP searches [ctascience2018, abdalla2021].
@@
For the baseline states of our sources, we adopt the best-fit spectral models reported in the 4th Fermi-LAT Source Catalog Data Release 4 (4FGL-DR4) [4fgl-dr3, 4fgl-dr4]. For Mrk 501 and PKS 2155-304, these are simple log-parabola functions. Since there are no observational constraints on the gamma-ray emission of blazars above a few tens of TeV, we conservatively assume that the validity of these models can be extended up to $10\,{\rm TeV}$ (thus not fully exploiting the CTAO energy range). In other works, an exponential cutoff factor around $1\,{\rm TeV}$ in rest-frame energy was added to these spectra (e.g. [abdalla2021]). We have checked that this modification does not significantly impact the extent of our final results.
@@
$\phi_{\rm obs}(E)=\phi_{\rm int}(E)e^{-\tau(E;z)}\,.$ (2.3)
Since the effect of EBL absorption is negligible in the Fermi-LAT energy range, we take the 4FGL models as representative of the intrinsic spectra of the sources, and apply the EBL absorption factor to account for flux reduction at higher energies. Several models of the EBL exist in the literature, both phenomenological and obtained from first principles, mostly providing consistent results [dominguez2011, franceschini2008, finke2010, gilmore2012, saldanalopez2021, deangelis2013]. In the following, we shall employ tabulated values of the optical depth as computed in the model of Domínguez et al. [dominguez2011]. We will discuss the impact of choosing different EBL models on our final results in Appendix D.
In addition to the Fermi-LAT baseline states, we consider for each source a flaring state taken from previous VHE measurements. Specifically, we take the Mrk 501 flare observed by the MAGIC-I telescope on 30 June 2005 [albert2007], which is again well-fit by a LP shape, and the historical PKS 2155-304 flare observed by H.E.S.S. on 29 July 2006 [aharonian2009], which follows an ECPL. These spectra are corrected for EBL absorption and, as in the previous case, we assume their validity up to $10\,{\rm TeV}$ and we extrapolate them down to $30\,{\rm GeV}$.
The selection of spectral parameters for each source and activity state is reported and discussed in greater detail in Appendix A.

### v2

# Projected sensitivity of CTAO to axion-like particles from blazars with a machine learning approach

@@ Abstract
Blazars are a class of active galactic nuclei, supermassive black holes located at the centres of distant galaxies characterised by strong emission across the entire electromagnetic spectrum, from radio waves to gamma rays. Their relativistic jets, closely aligned to the line of sight from Earth, are a rich and complex environment, characterised by the presence of strong magnetic fields over parsec-scale lengths. Owing to their cosmological distance from Earth, these sources serve as ideal targets to probe non-standard gamma-ray propagation. In particular, axion-like particles (ALPs) could be detected through their coupling to photons, which enables ALP-photon conversions in external magnetic fields, leading to distinct signatures in the blazars’ gamma-ray spectra. In this work, we estimate the potential of the Cherenkov Telescope Array Observatory (CTAO) to constrain the ALP parameter space by simulating observations of two bright blazars, Mrk 501 and PKS 2155-304. We obtain projected $2\sigma$ exclusion regions, demonstrating that CTAO will be able to consistently improve present limits thanks to its greater energy resolution and point-source sensitivity with respect to present ground-based gamma-ray telescopes. In addition to the standard statistical technique based on the likelihood ratio test, we further demonstrate the application of a new method based on machine learning classifiers, which may help in reducing the effect of systematic model-dependent uncertainties in future ALP searches.
@@ 1 Introduction
Axion-like particles (ALPs) are a kind of hypothetical, pseudo-scalar elementary particles whose existence is predicted by several extensions of the Standard Model. Their name is derived from the original axion model, which was formulated by Peccei and Quinn Peccei and Quinn (1977), Weinberg Weinberg (1978), and Wilczek Wilczek (1978) as a solution of the strong CP problem in the context of quantum chromodynamics (QCD). While the existence of this original axion was soon ruled out Donnelly et al. (1978); Zehnder (1981), several “invisible axion” models were proposed, such as KSVZ Kim (1979); Shifman et al. (1980) and DFSZ Dine et al. (1981); Zhitnitsky (1980), which could elude experimental constraints while still solving the strong CP problem. In addition to these, it became clear that axion-like particles could be a more general prediction of beyond-the-Standard-Model theories such as Grand Unification Theories and string theories. These ALPs are not necessarily related to the original Peccei-Quinn model, but, like the QCD axion, they often arise as pseudo-Nambu-Goldstone bosons of spontaneously broken symmetry groups. Axions and ALPs are currently of great interest in cosmology and astrophysics, as they are among the best-motivated candidates as constituents of dark matter. Reviews of models and experimental searches for both QCD axions and ALPs can be found e.g. in refs. Jaeckel and Ringwald (2010); Ringwald (2012); Irastorza and Redondo (2018); Di Luzio et al. (2020).
@@
The most important feature for the experimental detection of ALPs is their effective two-vertex coupling to photons, whose strength is controlled by the coupling constant $g_{a\gamma}$. Such a coupling is the basis for the most common approach to ALP searches, which exploits the expected ALP-photon mixing in the presence of external magnetic fields Sikivie (1983); Raffelt and Stodolsky (1988). This effect is employed, for example, in beam regeneration (or “light shining through walls”, LSW) experiments, such as ALPS-II Bähre et al. (2013), where a laser beam is shot against an opaque screen; in the presence of a strong enough magnetic field, some photons in the beam should be able to convert into ALPs, pass through the screen and oscillate back to photons on the other side. Other experiments for the direct detection of ALPs include helioscopes, like CAST Zioutas et al. (1999) and the foreseen IAXO Armengaud et al. (2019), and resonant cavity experiments such as ADMX Du et al. (2018) and MADMAX Brun et al. (2019).
@@
As a general property of QCD axions, the axion-photon coupling $g_{a\gamma}$ is proportional to the axion mass $m_{a}$, as both quantities are inversely proportional to the Peccei-Quinn energy scale $f_{a}$ (e.g. Irastorza and Redondo (2018)). On the other hand, in the more general case of ALPs there is typically no fixed relationship between $m_{a}$ and $g_{a\gamma}$, so that they are independent coordinates spanning a 2-dimensional parameter space.
@@
Astrophysical gamma-ray sources characterised by one or more magnetic field environments along their line of sight to Earth offer a natural framework to probe ALP-related effects, which has been extensively discussed in literature Hooper and Serpico (2007); Mirizzi et al. (2007, 2008); De Angelis et al. (2007, 2008, 2011); Horns et al. (2012); Batković et al. (2021); Galanti and Roncadelli (2022). Gamma-ray polarization effects could in principle observed, as ALPs only couple to the photon state parallel to the external transverse magnetic field; however, such effects are not detectable with present-day instruments Bassan et al. (2010); Galanti (2024). On the other hand, two main measurable spectral features may be investigated, whose extent generally depends on the field strength and topology, ALP parameters, source distance, and energy range of interest. The first is the presence of energy-dependent wiggles superimposed on the spectral shape, due to the peculiar shape of the photon survival probability $P_{\gamma\gamma}$ as a function of energy. The second is a spectral hardening at very high energy (VHE), due to gamma-ray photons converting into ALPs in the magnetic field of the source and back into photons in the Milky Way. This mechanism would allow the ALP-photon beam to travel cosmological distances, eluding the absorption that is normally experienced by VHE gamma rays due to the pair production (Breit-Wheeler scattering Breit and Wheeler (1934)) on the soft photons making up the extragalactic background light (EBL) Domínguez et al. (2011); Franceschini et al. (2008); Finke et al. (2010); Gilmore et al. (2012); Saldana-Lopez et al. (2021); De Angelis et al. (2013). The resulting effect, analogous to a “cosmic LSW experiment”, is expected to be more significant in the spectral tails of distant sources, for which the EBL absorption in the absence of ALPs would be strongest.
@@
A typical target of these studies is represented by blazars. These sources are a particular class of active galactic nuclei (AGN), supermassive black holes characterised by accretion disks and strong relativistic jets at a small angle ($\lesssim 20\degree$) with the line of sight from Earth Padovani et al. (2017). Their broadband spectral energy distribution (SED) is characterised by a low-energy component due to the synchrotron radiation emitted by accelerated electrons in the jet, peaked between infrared and X-ray frequencies, and higher-energy emission in the gamma rays that may be produced either by leptonic or hadronic mechanisms Böttcher et al. (2013); Cerruti (2020). Leptonic mechanisms typically assume the gamma-ray emission to be originated by the same electron population that produces the synchrotron radiation via inverse Compton scattering, either on the synchrotron radiation field itself (synchrotron self-Compton, SSC) or on an external low-energy photon field (such as thermal photons from the broad line region, accretion disk or dusty torus, or the cosmic microwave background). In hadronic models, high-energy photons are emitted by a population of relativistic protons and other nuclei in the jet, typically via synchrotron radiation or photo-meson production, with the associated production of neutrinos. Based on their optical properties, blazars can be classified between BL Lacertae (BL Lac) objects and flat spectrum radio quasars (FSRQs). In addition, they can be distinguished between high-, intermediate- and low-synchrotron peaked (HSP, ISP and LSP) depending on their peak synchrotron frequency. Many blazars display significant spectral variability over timescales ranging from years down to minutes Böttcher (2019), with high-flux flaring states representing important opportunities of detection or even discovery of new sources. Despite being a very rare class of objects, almost all extragalactic sources detected in gamma-rays are associated with blazars.
@@
The relativistic jets of blazars typically feature magnetic fields of order $\mathcal{O}(\rm mG)$ to $\mathcal{O}(\rm G)$ over $\mathcal{O}(\rm pc)$ scales. These objects may also be located inside galactic clusters with complex magnetic field structures of $\mathcal{O}(\rm\upmu G)$ intensity over $\mathcal{O}(\rm Mpc)$ scales; a notable example of this environment is represented by the Perseus galaxy cluster, which hosts the Seyfert galaxy NGC 1275. Additionally, being extragalactic sources, blazars are located at a cosmological distance from Earth (represented by the redshift parameter $z$), which makes EBL absorption effects important. All these characteristics may allow for hints of ALP-photon conversions in the gamma-ray spectra of blazars. Several constraints on the ALP parameter space have been obtained by analysing gamma-ray data from different blazars Abramowski et al. (2013); Jacobsen et al. (2023); Davies et al. (2023) and from the Perseus galaxy cluster Ajello et al. (2016); Abe et al. (2024a).
@@
In this context, the next-generation Cherenkov Telescope Array Observatory (CTAO [^1] https://www.ctao.org/, formerly known as CTA) will be an ideal instrument to probe this kind of phenomena. The observatory will be made up of three different types of imaging atmospheric Cherenkov telescopes (IACTs), distributed in arrays over two sites in the Northern and Southern hemispheres Acharya et al. (2013). The Northern Array will be located at Observatorio del Roque de los Muchachos in La Palma, Canary Islands, and will be made up of 4 Large-Sized Telescopes (LSTs) and 9 Medium-Sized Telescopes (MSTs); the Southern Array, in Paranal, Chile, will feature 14 MSTs and 37 Small-Sized Telescopes (SSTs). These different configurations were designed to obtain the best sensitivity to sources visible from the two hemispheres, with the Southern array optimised to detect the highest energy photons (up to several hundreds of TeV) from the Galactic centre. In addition to this baseline “Alpha” configuration, an upgrade is foreseen by the Italian CTA+ Program [^2] https://pnrr.inaf.it/progetto-ctaplus/ to add 2 LSTs and 5 more SSTs to the Southern site Loporchio (2025). An alternative configuration has also been proposed, using telescopes of the Schwarzschild-Couder type Adams et al. (2021) to complement the standard MSTs (which feature a Davies-Cotton optical design). CTAO will obtain a significant improvement in all areas of performance with respect to present-days IACT systems: in particular, its greatly enhanced energy resolution and point-source sensitivity will make it an ideal instrument to study the spectra of distant sources in search of potential non-standard spectral features Acharya et al. (2018). The CTAO sensitivity over the ALP parameter space has been estimated by simulating observations of various blazars Meyer et al. (2014); Meyer and Conrad (2014) and of the Perseus galaxy cluster Abdalla et al. (2021).
@@
Almost all of the previously cited works rely on the definition of some kind of test statistic (TS) based on the logarithm of the likelihood ratio (equivalently, the difference in log-likelihoods) to test the ALP hypothesis on real or simulated data Jacobsen et al. (2023); Davies et al. (2023); Ajello et al. (2016); Abe et al. (2024a); Meyer et al. (2014); Meyer and Conrad (2014); Abdalla et al. (2021). In mathematical terms, the TS is given by
${\rm TS}=-2\ln\frac{\mathcal{L}(H_{1})}{\mathcal{L}(H_{0})}\,$, (1)
@@ 2.1 Source selection
The observation of AGNs is one of the Key Science Projects (KSPs) proposed for CTAO, to be carried out within the first 10 years of operation of the observatory Acharya et al. (2018). The programme of this KSP will be organized in three parts: long-term monitoring of known AGNs, search and follow-up of flares, and measurement of high-quality spectra for different AGN classes and redshifts. One of the declared goals of the AGN KSP is to study possible imprints of new fundamental physics on the cosmological propagation of gamma-rays, including ALP searches Acharya et al. (2018); Abdalla et al. (2021).
@@
For the baseline states of our sources, we adopt the best-fit spectral models reported in the 4th Fermi-LAT Source Catalog Data Release 4 (4FGL-DR4) Abdollahi et al. (2022); Ballet et al. (2024). For Mrk 501 and PKS 2155-304, these are simple log-parabola functions. Since there are no observational constraints on the gamma-ray emission of blazars above a few tens of TeV, we conservatively assume that the validity of these models can be extended up to $10\,{\rm TeV}$ (thus not fully exploiting the CTAO energy range). In other works, an exponential cutoff factor around $1\,{\rm TeV}$ in rest-frame energy was added to these spectra (e.g. Abdalla et al. (2021)). We have checked that this modification does not significantly impact the extent of our final results.
@@
$\phi_{\rm obs}(E)=\phi_{\rm int}(E)e^{-\tau(E;z)}\,.$ (4)
Since the effect of EBL absorption is negligible in the Fermi-LAT energy range, we take the 4FGL models as representative of the intrinsic spectra of the sources, and apply the EBL absorption factor to account for flux reduction at higher energies. Several models of the EBL exist in the literature, both phenomenological and obtained from first principles, mostly providing consistent results Domínguez et al. (2011); Franceschini et al. (2008); Finke et al. (2010); Gilmore et al. (2012); Saldana-Lopez et al. (2021); De Angelis et al. (2013). In the following, we shall employ tabulated values of the optical depth as computed in the model of Domínguez et al. Domínguez et al. (2011). We discuss the impact of choosing different EBL models on our final results in Appendix A.
In addition to the Fermi-LAT baseline states, we consider for each source a flaring state taken from previous VHE measurements. Specifically, we take the Mrk 501 flare observed by the MAGIC-I telescope on 30 June 2005 Albert et al. (2007), which is again well-fit by a LP shape, and the historical PKS 2155-304 flare observed by H.E.S.S. on 29 July 2006 Aharonian et al. (2009), which follows an ECPL. These spectra are corrected for EBL absorption and, as in the previous case, we assume their validity up to $10\,{\rm TeV}$ and we extrapolate them down to $30\,{\rm GeV}$.
The selection of spectral parameters for each source and activity state is reported and discussed in greater detail in Appendix B.